# Utilizing smartphone sensors for accurate solar irradiance measurement and educational purposes

### v2

@@ Abstract
The global transition towards cleaner and more sustainable energy production is a major challenge. We present an innovative solution by utilizing smartphone light sensors to measure direct normal solar irradiance, the primary component of ground-level solar radiation. We provide comprehensive guidelines for calibrating the sensor using two methods: a professional reference measurement and clear-sky satellite estimates. The latter method is particularly advantageous in resource-constrained environments. Once calibrated, the smartphone becomes a valuable tool for measuring the solar resource. After calibrating the device, we propose an instructional laboratory focusing on the physics of solar radiation and its interaction with the Earth’s atmosphere, exploring solar variations across locations, cloud conditions, and time scales. By integrating irradiance values measured throughout a day the daily irradiation can be estimated. This approach enhances students’ understanding of solar radiation attenuation and its relationship with atmospheric interactions. This method offers a practical and educational solution for promoting renewable energy knowledge and addressing the challenges of the energy transition.
@@ 1 Introduction
The rise of renewable energy sources, particularly solar and wind power, is an unstoppable force [1, 2]. This transformative shift in energy production and consumption necessitates a workforce equipped with specialized knowledge and skills. However, the incorporation of renewable energy topics in science and engineering curricula, especially in physics courses, has often been marginalized. Specifically, the study of solar radiation is predominantly approached from a theoretical standpoint, potentially due to the technical complexities and high costs associated with measurement instruments. As a result, experiments involving the measurement of solar irradiance and the determination of its uncertainty are typically reserved for advanced programs or specialized laboratories.
The importance of the contents associated with solar energy transcends its specific aspects. In fact, this subject matter provides a platform to develop various cross-cutting competencies, including measurement techniques and environmental stewardship. In summary, the dissemination of instructional laboratories focusing on clean and environmentally sustainable methods of energy production serves the dual purpose of encouraging responsible citizen engagement in public policy discussions and training professionals who can contribute to the growth of the sector.
The proposal of this article makes it possible to break down the cost barrier making it possible to incorporate in undergraduate experimental physics courses. The cost barrier is eliminated through the use of smartphones as a measurement instrument, which have also shown innovative contributions to the teaching of Physics for different thematic areas [3, 4, 5]. We show here how to use a smartphone’s light sensor [6] to measure broadband direct solar irradiance at normal incidence (DNI), the main component of solar radiation at ground level. Direct irradiance is the portion of the incident radiation that arrives directly from the solar disk without being absorbed or scattered in the Earth’s atmosphere. This component is essential to evaluate the performance of concentrating solar applications and to estimate the solar irradiance available on an inclined plane, necessary for solar photovoltaic and low-temperature solar thermal applications in which it is usual practice to tilt the solar collection surfaces. Here, we describe two procedures for calibrating the sensor, one based on a professional reference measurement and the other using publicly-available satellite estimates of ground level clear-sky irradiance. In both cases we pay special attention to non-trivial implementation details. We compare the calibration procedures and show that both are feasible. By bridging this gap about calibration, we enable the measurements and instructional laboratory to be done with only the smartphone and manual positioning, if required. The instructional laboratory proposed here has a focus on solar resource fundamentals, which are required to understand any form of solar energy application.
@@
This article is organized as follows. Next section introduces the basic aspects of solar radiation modeling, the various magnitudes and usual geometric calculations in the area of solar energy and lighting. Section 3 shows the experimental setup carried out in this work for the professional calibration of the illuminance sensor of a smartphone for the measurement of DNI. This section also introduces the solar satellite estimates, which are used here as an alternative reference data set for smartphone calibration. These data are publicly available for download on dedicated websites with global coverage, so their use does not represent a limitation. Section 4 describes the calibration process and its uncertainty evaluation against both sets of data. Once the sensors are calibrated it is possible to propose different activities. Specifically in Section 5 we propose several activities that can conform an instructional laboratory in this area. Finally, Section 6 summarizes the conclusions.
@@ 2 Theoretical framework: solar radiation
The solar irradiance, , is the incident power per unit normal surface whose beam comes from the Sun. The solar irradiance at the top of the atmosphere, , varies because of two factors: a variation of about % due to the elliptical character of the Earth’s orbit around the Sun, and also, it present small variations due to oscillations in the solar activity, typically below % [7]. The averaged solar irradiance at the top of the atmosphere on a surface normal to the Earth-Sun direction and when the Earth is at a distance equal to the mean Earth-Sun distance (1 Astronomical Unit or AU) is known as the solar constant [8], Wm. In this way, the seasonal variation of is obtained by multiplying the solar constant by the orbital factor, , accounting for the variation of the Earth-Sun distance and it can be approximated with an uncertainty of 0.25% by
@@
where is the ordinal day number (going from 1, Jan. 1st. to 365 Dec. 31) [9, 10]. The extraterrestrial irradiance at normal incidence is obtained then as .
@@
Once solar irradiance penetrates the Earth’s atmosphere, it interacts with various atmospheric components such as air, aerosols, water vapor, and cloudiness. This interaction leads to scattering in multiple directions, with some of the irradiance being absorbed by these components, while the remaining portion is reflected back into space. A certain percentage of the irradiance originating from the top of the atmosphere eventually reaches the Earth’s surface. It does so through two distinct paths: (i) directly from the solar disk, without undergoing scattering or absorption by atmospheric components and (ii) from the rest of the sky, following a series of complex processes involving multiple scattering. The combination of these two components on a horizontal plane is known as the global horizontal irradiance (GHI), denoted as , which represents the solar energy magnitude most commonly measured on the Earth’s surface. Several methods can be employed to measure this quantity, including the use of photodiodes, calibrated photovoltaic cells, or thermopile pyranometers. Among these options, the last offers the highest precision.
@@
Direct normal irradiance, denoted as , is less frequently measured since its continuous measurement requires fine solar tracking mechanisms that ensure the measuring equipment to be aligned at all times pointing to the solar disk. The measuring instrument, the pyroheliometer, is equipped with a collimating tube that filters any irradiance that does not come from its normal direction, with a convention aperture of 5 sr of solid angle, corresponding to a typical solar disk. The size of the solar disk observed from Earth depends on atmospheric conditions. In the presence of high humidity, for example, the perceived solar disk enlarges due to a larger size of the circumsolar region. The solid angle of 5 sr associated with the solar disk is, in effect, a convention value. The standard that classifies solar radiation measurement instruments is the ISO 9060:2018 which establishes categories according to the quality of the equipment (offset, angular error, response time, among others) and the corresponding uncertainty. After measuring the DNI, the atmospheric transmittance can be estimated as , a quantity that allows quantifying the attenuation of the irradiance in the atmosphere, in particular, in the presence of cloudiness.
@@
In the upcoming section our focus is on presenting a straightforward clear-sky model that relies on the Lambert-Beer-Bouger law. To establish this model, we must first define several geometric quantities. One quantity is the solar zenith angle, denoted as , which represents the angle between the direction of the Sun and the local vertical (referred to as the local zenith) as shown in Fig. 1a. The cosine of this angle appears recurrently in expressions related to solar radiation, especially for magnitudes projected onto the horizontal plane, and its calculation is carried out according to [10]
@@
where is the latitude, is the solar declination angle and is the hour angle shown in Fig. 1b. Latitude is the angle between the Earth’s equator (parallel 0°) and the site of interest (indicated by in the figure), along the observer’s meridian. By convention, latitudes are positive north of the equator and negative south of the equator. Solar declination is the angle formed by the Earth-Sun line with the Earth’s equatorial plane, and can be calculated in radians with good precision through the expression:
@@
Finally, the hour angle is the angle on the equatorial plane between the observer’s meridian and the solar meridian. This angle varies with the apparent position of the Sun respect to the Earth and it is calculated from the time label associated with each measurement. Figure 1b also depicts a fourth relevant angle, , the longitude of the observer measured from the Greenwich meridian (°).
@@
Indeed, this angle is at solar noon ( hours), i.e. when the solar meridian coincides with the observer’s meridian, and grows at a rate of radians per hour, the speed of rotation of the Earth. To complete the calculation, it remains to link the solar local time with the local standard time, , expressed according to a given UTC time zone associated to a central meridian, . For example, the time in UTC-3 is associated with a meridian of °. The relationship between both hours includes the so-called equation of time, , and is defined by
where is the local standard time expressed in hours and fractions, and are the signed longitude in decimal degrees (negative for West longitudes and positive for East longitudes) of the site and the reference UTC, respectively, and is expressed in minutes. Based on Spencer’s developments[11], can be calculated as a Fourier expansion in terms of the variable as
2.1 Air mass
The mass of air or relative optical mass, , is a dimensionless quantity that is defined as the quotient between the amount of mass of a certain component -th of the atmosphere that a beam of radiation covers in its trajectory and the one that it would cover in a vertical path, that is, in the direction of the zenith. If a non-anisotropic flat atmosphere is assumed for its calculation, the following simplified and purely geometric expression is obtained [9, 10]:
The uncertainty associated with this expression due to neglecting the terrestrial curvature and the refraction phenomena grows as the zenith angle is larger. However, the expression presents an uncertainty of about % for ° [9], and it is adequate for zenith angles between 0° and 70°. There were proposed more precise expressions that can be used for large zenith angles, around 80-90°, such as that of Kasten and Young [12]. Here, as we do not consider measurements very early in the morning or very late at sunset the previous expression, Eq. 7, results appropriate.
2.2 Lambert-Beer-Bouger law
@@
where is the direct spectral irradiance, is the spectral extraterrestrial irradiance corrected by the orbital factor, is the optical depth of the atmosphere and is the air mass. This equation can be derived from the differential version of the Lambert-Beer-Bouger modeling the atmosphere by a set of layers , so as the transmissivity can be expressed as . The total transmissivity results from the product of the layers, and therefore includes the effect of all different components. This is a regular assumption when modeling the interaction between the Sun’s radiation and the atmosphere [13, 14].
Clear-sky models adopt the Lambert-Beer-Bouger law to describe the direct normal irradiance under ideal atmospheric conditions. In these conditions, the attenuating components encompass various factors such as air molecules (O, N, Ar), which are responsible for Rayleigh scattering, as well as water vapor, aerosols, ozone, and other minor gases. Ozone, although crucial for life on Earth due to its role in attenuating ultraviolet radiation, has a relatively minor contribution across the entire solar spectrum. A Rayleigh atmosphere refers to a pristine and dry atmospheric state where only the attenuation mechanism of molecular scattering is at play, leading to a clean and transparent atmosphere.
Numerous clear-sky models have been developed based on these concepts [15]. Among them, the ESRA (European Solar Radiation Atlas) model [16] strikes a favorable balance between simplicity and precision, making it suitable for implementation within the framework of an university experimental practice. This model operates using the concept of global optical depth, denoted as , which encompasses the entire solar spectrum. By incorporating the global optical depth of Rayleigh extinction, denoted as , we can express , where represents Linke’s turbidity which quantifies the number of clean and dry atmospheres that would need to be stacked to achieve the same level of attenuation observed in the real atmosphere. Consequently, by adjusting a single parameter, , based on ground measurements it becomes feasible to construct a straightforward model for estimating DNI under clear-sky conditions as . There are several ways to approximate the Rayleigh optical depth [17, 18, 12, 19, 20]. Here we use the formulation of Kasten [20], used in the ESRA model:
@@ 2.3 Illuminance
Photometry is the area of knowledge that is responsible for measuring the light perceived by the human eye. This quantity depends on the sensitivity of the human eye to different wavelengths in the visible region of the electromagnetic spectrum. Each wavelength has its relative weight in the response of the human eye depending on the lighting conditions (good or poor) in which the observer is. In typical lighting conditions, corresponding to a real situation at Sun, it is possible to relate the illuminance, , measured in lumens per unit area, lm/m or lx, with the spectral irradiance according to
@@
Consequently, establishing a precise relationship between illuminance, denoted as , and broadband solar irradiance is not a straightforward task, as it depends on the spectral composition of solar irradiance at ground level within a specific portion of the spectrum. This composition, in turn, is influenced by the atmospheric conditions. This scenario resembles the calibration process for photovoltaic radiometers used to measure solar irradiance. These devices possess distinct spectral responses across different regions of the solar spectrum and are calibrated by comparing them to pyranometric radiometers with a flat spectral response (broadband) encompassing the entire solar spectrum. As a first approximation, these spectral differences can be disregarded, and the customary approach involves employing a constant or global calibration curve, determined under clear-sky conditions, to account for these effects [21]. This calibration methodology is adopted in the present study.

### v3

@@ Abstract
The use of solar energy has experienced explosive growth in recent years due to the comparative advantages it offers. Here, we present an innovative tool for using the light sensor of smartphones for solar irradiance experiments in particular to measure the primary component of ground-level solar radiation: the direct normal solar irradiance. We provide comprehensive guidelines for using this sensor in a variety of situations and show how to calibrate it using easy available clear-sky satellite data. In addition, as an alternative approach, we also incorporate a comparison with professional reference measurement estimates to demonstrate the consistency of the proposed methods. This experiment can be performed with the vast majority of available smartphones, some readily available materials plus an open location in clear sky conditions for a lapse of a couple of hours. In this way, the smartphone becomes a valuable tool for measuring the solar resource and to propose several experiments focusing on the physics of solar radiation and its interaction with the Earth’s atmosphere, exploring solar variations across locations, cloud conditions, and time scales.
keywords:
[…]
@@ 1 Introduction
Photovoltaic energy has experienced spectacular growth in recent years BP (2021); REN21 (2021). This growth contrasts with the little attention given to solar resources in science and engineering curricula. In general, the study of solar radiation is predominantly approached from a theoretical standpoint, potentially due to the technical complexities and high costs associated with measurement instruments. As a result, experiments involving the measurement of solar irradiance and the determination of its uncertainty are typically reserved for advanced programs or specialized laboratories. It is also worth mentioning that this subject matter provides a platform to develop various cross-cutting competencies, including measurement techniques and environmental stewardship.
The main objective of this work is to show how to calibrate a light sensor like the ones incorporated in most smartphones to measure the solar resource. The achievement of this objective allows to dispose of an instrument to measure solar radiation for other activities and/or laboratories and also provides the opportunity to learn about the solar resource in general. In particular, we show how to measure broadband direct solar irradiance at normal incidence (DNI), the main component of solar radiation at ground level. Direct irradiance is the portion of the incident radiation that arrives directly from the solar disk without being absorbed or scattered in the Earth’s atmosphere. This component is essential to evaluate the performance of concentrating solar applications and to estimate the solar irradiance available on an inclined plane, necessary for solar photovoltaic and low-temperature solar thermal applications in which it is usual practice to tilt the solar collection surfaces.
The use of smartphones as a measurement instrument, which have also shown innovative contributions to the teaching of Physics due to the availability of various built-in sensors (for recent reviews see Ref. Monteiro and Martí (2022); Gil and Di Laccio (2017)). Although the type of sensor varies the type of sensor varies considerably depending on the specific model, in general, acceleration, magnetic field or luminosity sensors are the most common. The latter, which we will use in this work, has been used in relatively few experiments. It is worth mentioning its use for experimentation with polarized waves Monteiro et al. (2017) or in problems motivated in astronomical situations Di-Laccio et al. (2021). Among the advantages of using smartphone sensors we can mention that it is possible to measure simultaneously with several sensors mon (2019) or supplementted with video analysis mon (2021). In the framework of experiences aimed at students, the determination of uncertainty and the study of the fluctuations of these sensors cannot be left aside Monteiro et al. (2021)
To reach our goal, we describe two procedures for calibrating the sensor, one based on a professional reference measurement and the other using publicly-available satellite estimates of ground level clear-sky irradiance. We compare the calibration procedures and show that both are feasible. By bridging this gap about calibration, we enable the measurements and instructional laboratory to be done with only the smartphone and manual positioning, if required. We must emphasize that the professional equipment mentioned in the paper, including tracking, is not indispensable to carry out the proposal, and is included only to show the reader that the smartphone is a calibrable and reliable instrument.
@@
As the topics concernig the solar resource are not usually covered in physics courses, next section introduces the various magnitudes and usual geometric calculations in the area of solar energy and lighting. Section 3 shows the experimental setup carried out in this work for the calibration of the illuminance sensor of a smartphone for the measurement of DNI. This section also introduces the solar satellite estimates, which are used here as an alternative reference data set for smartphone calibration. Section 4 describes the calibration process and its uncertainty evaluation. Finally, Section 5 summarizes the conclusions.
@@ 2 Theoretical framework: solar radiation
2.1 Basis of the interaction between solar radiation and the atmosphere
@@
The solar irradiance, , is the incident power per unit normal surface whose beam comes from the Sun. The solar irradiance at the top of the atmosphere, , varies because of two factors: a variation of about % due to the elliptical character of the Earth’s orbit around the Sun, and also, it present small variations due to oscillations in the solar activity, typically below % Fröhlich (2006). The averaged solar irradiance at the top of the atmosphere on a surface normal to the Earth-Sun direction and when the Earth is at a distance equal to the mean Earth-Sun distance (1 Astronomical Unit or AU) is known as the solar constant Kopp and Lean (2011), Wm-2. In this way, the seasonal variation of is obtained by multiplying the solar constant by the orbital factor, , accounting for the variation of the Earth-Sun distance and it can be approximated with an uncertainty of 0.25% by
@@
where is the ordinal day number (going from 1, Jan. 1st. to 365 Dec. 31) Iqbal (1983); Duffie and Beckman (2013). The extraterrestrial irradiance at normal incidence is obtained then as .
@@
Once solar irradiance penetrates the Earth’s atmosphere, it interacts with various atmospheric components such as air, aerosols, water vapor, and cloudiness. This interaction leads to scattering in multiple directions, with some of the irradiance being absorbed by these components, while the remaining portion is reflected back into space. The combination of these two components on a horizontal plane is known as the global horizontal irradiance (GHI), denoted as , which represents the solar energy magnitude most commonly measured on the Earth’s surface. Several methods can be employed to measure this quantity, including the use of photodiodes, calibrated photovoltaic cells, or thermopile pyranometers. Among these options, the last offers the highest precision.
@@
Direct normal irradiance, denoted as , is less frequently measured since its continuous measurement requires fine solar tracking mechanisms that ensure the measuring equipment to be aligned at all times pointing to the solar disk. The measuring instrument, the pyroheliometer, is equipped with a collimating tube that filters any irradiance that does not come from its normal direction, with a convention aperture of 5 sr of solid angle, corresponding to a typical solar disk. The size of the solar disk observed from Earth depends on atmospheric conditions. In the presence of high humidity, for example, the perceived solar disk enlarges due to a larger size of the circumsolar region. The solid angle of 5 sr associated with the solar disk is, in effect, a convention value. The standard that classifies solar radiation measurement instruments is the ISO 9060:2018 which establishes categories according to the quality of the equipment (offset, angular error, response time, among others) and the corresponding uncertainty. After measuring the DNI, the atmospheric transmittance can be estimated as , a quantity that enables understanding the amount of available solar radiation for capture and conversion into solar energy.
@@
It is useful to define the main angles related to solar radiation. One of them is the solar zenith angle, denoted as , which represents the angle between the direction of the Sun and the local vertical (referred to as the local zenith) as shown in Fig. 1. The cosine of this angle appears recurrently in expressions related to solar radiation, especially for magnitudes projected onto the horizontal plane, and its calculation is carried out according to Duffie and Beckman (2013)
@@
where is the latitude, is the solar declination angle and is the hour angle shown in Fig. 1. Latitude is the angle between the Earth’s equator (parallel 0°) and the site of interest (indicated by in the figure), along the observer’s meridian. By convention, latitudes are positive north of the equator and negative south of the equator. Solar declination is the angle formed by the Earth-Sun line with the Earth’s equatorial plane, and can be calculated in radians with good precision through the expression:
@@
Finally, the hour angle is the angle on the equatorial plane between the observer’s meridian and the solar meridian. This angle varies with the apparent position of the Sun respect to the Earth and it is calculated from the time label associated with each measurement. Figure 1 also depicts a fourth relevant angle, , the longitude of the observer measured from the Greenwich meridian (defined as °).
@@
Indeed, this angle vanishes at solar noon ( hours), i.e. when the solar meridian coincides with the observer’s meridian, and grows at a rate of radians per hour, the speed of rotation of the Earth. To complete the calculation, it remains to link the solar local time with the local standard time, , expressed according to a given UTC time zone associated to a central meridian, . For example, the time in UTC-3 is associated with a meridian of °. The relationship between both hours includes the so-called equation of time, , and is defined by
where is the local standard time expressed in hours and fractions, and are the signed longitude in decimal degrees (negative for West longitudes and positive for East longitudes) of the site and the reference UTC, respectively, and is expressed in minutes. can be calculated as a Fourier expansion as proposed by Spencer Spencer (1971).
2.2 Attenuation in the atmosphere
The mass of air or relative optical mass, , is a dimensionless quantity that is defined as the quotient between the amount of mass of a certain component -th of the atmosphere that a beam of radiation covers in its trajectory and the one that it would cover in a vertical path, that is, in the direction of the zenith. Under the hypothesis of a non-anisotropic flat atmosphere, we can assume Iqbal (1983); Duffie and Beckman (2013). The uncertainty associated with this expression due to neglecting the terrestrial curvature and the refraction phenomena grows as the zenith angle is larger. However, the expression presents an uncertainty of about % for ° Iqbal (1983), and it is adequate for zenith angles between 0° and 70°. There were proposed more precise expressions that can be used for large zenith angles, around 80-90°, such as that of Kasten and Young Kasten and Young (1989). Here, as we do not consider measurements very early in the morning or very late at sunset the previous expression, the previous approximation results appropriate.
@@ 2.2 Attenuation in the atmosphere
where is the direct spectral irradiance, is the spectral extraterrestrial irradiance corrected by the orbital factor, is the optical depth of the atmosphere and is the air mass. This equation can be derived from the differential version of the Lambert-Beer-Bouger modeling the atmosphere by a set of layers , so as the transmissivity can be expressed as . The total transmissivity results from the product of the layers, and therefore includes the effect of all different components. This is a regular assumption when modeling the interaction between the Sun’s radiation and the atmosphere Gueymard (2008); Oumbe et al. (2014).
Clear-sky models adopt the Lambert-Beer-Bouger law to describe the direct normal irradiance under ideal atmospheric conditions. In these conditions, the attenuating components encompass various factors such as air molecules (O2, N2, Ar), which are responsible for Rayleigh scattering, as well as water vapor, aerosols, ozone, and other minor gases. Ozone, although crucial for life on Earth due to its role in attenuating ultraviolet radiation, has a relatively minor contribution across the entire solar spectrum. A Rayleigh atmosphere refers to a pristine and dry atmospheric state where only the attenuation mechanism of molecular scattering is at play, leading to a clean and transparent atmosphere.
Numerous clear-sky models have been developed based on these concepts Gueymard (2012). Among them, the ESRA (European Solar Radiation Atlas) model Rigollier et al. (2000) strikes a favorable balance between simplicity and precision, making it suitable for implementation within the framework of an university experimental laboratory. This model operates using the concept of global optical depth, denoted as , which encompasses the entire solar spectrum. By incorporating the global optical depth of Rayleigh extinction, denoted as , we can express , where represents Linke’s turbidity which quantifies the number of clean and dry atmospheres that would need to be stacked to achieve the level of attenuation observed in the real atmosphere. Consequently, by adjusting a single parameter, , based on ground measurements it becomes feasible to construct a simple model for estimating DNI under clear-sky conditions as . Among the several ways to approximate the Rayleigh optical depth Kasten (1984); Louche et al. (1986); Kasten and Young (1989); Grenier et al. (1994); Kasten (1996), here we use the Kasten’s formulation Kasten (1996).
@@ 2.3 Illuminance
Photometry is the area of knowledge that is responsible for measuring the light perceived by the human eye. This quantity depends on the sensitivity of the human eye to different wavelengths in the visible region of the electromagnetic spectrum. Each wavelength has its relative weight in the response of the human eye depending on the lighting conditions (good or poor) in which the observer is. In typical lighting conditions, corresponding to a real situation at Sun, it is possible to relate the illuminance, , measured in lumens per unit area, lm/m2 or lx, with the spectral irradiance according to
@@
Consequently, establishing a precise relationship between illuminance, denoted as , and broadband solar irradiance is not a straightforward task, as it depends on the spectral composition of solar irradiance at ground level within a specific portion of the spectrum. This composition, in turn, is influenced by the atmospheric conditions. This scenario resembles the calibration process for photovoltaic radiometers used to measure solar irradiance. These devices possess distinct spectral responses across different regions of the solar spectrum and are calibrated by comparing them to pyranometric radiometers with a flat spectral response (broadband) encompassing the entire solar spectrum. As a first approximation, these spectral differences can be disregarded, and the customary approach involves employing a constant or global calibration curve, determined under clear-sky conditions, to account for these effects ISO 9847:1992 (1992). This calibration methodology is adopted in the present study.